The sum of a number and –16 is 3. Find the number.
Found 2 solutions by neta, bucky:
Answer by neta(14)

first you need to know what you are looking for i.e the missing nos
so let x represent the missing number
therefore
x+-(16)=3
x-16=3
take 16 ove and the negative sign changes
x=3+16
x=19

Answer by bucky(2189)

Let x be the unknown number. Then the sum of it and -16 is represented by:
.
x + (-16)
.
And the problem tells you that this is to equal 3. In equation form this can be written as:
.
x + (-16) = 3
.
Since the parentheses are preceded by a plus sign, they can be removed without changing
the sign of the term inside the parentheses. When you do that removal, the equation
becomes:
.
x - 16 = 3
.
to solve for x get rid of the -16 on the left side by adding 16 to both sides. The addition
of 16 to both sides results in:
.
x - 16 + 16 = 3 + 16
.
On the left side the -16 and + 16 cancel each other out. And on the right side the
3 + 16 is 19. So the equation becomes:
.
x = 19
.
The missing number that solves this problem is +19
.
Hope this helps you to understand the basics of this problem and shows you a way to logically
solve it.